A company purchased a tractor on January 1, 2021, for $65,000. The tractor's useful life is estimated to be 30,000 miles with an expected residual value of $5,000. If the company used the tractor 5,000 miles in 2021 and 3,000 miles in 2022, what is the balance for accumulated depreciation at the end of 2022 using the activity-based method? Multiple Choice $6,000. $16,000. $38,000. $10,000.
Answer #1

Annual depreciation=(Cost-Residual value)/Useful Life

=(65,000-5,000)/30,000

=$2/mile

Hence depreciation for:

2021=(2*5000)=$10,000

2022=(2*3000)=$6,000

Accumlated depreciation=(10,000+6000)=$16,000